OpenStudy (anonymous):

Which answer shows this equation written in standard form? 8 – 1(x – y) = –3x + 5 A. 2x + y = –3 B. 2x – y = –3 C. –4x + y = –3 D. 2x + y = 13

OpenStudy (anonymous):

plz help.fan and medal

OpenStudy (anonymous):

@mathmate

OpenStudy (anonymous):

@peterg2001

OpenStudy (anonymous):

@bibby

OpenStudy (mathmate):

Hint: Standard form is in the form such as 5x+3y=23 with unknowns on the left, and constants on the right. To convert an equation to standard form, distribute any parentheses, and swap unknowns to the left, and numbers to the right. ------------------------- for \(example\) 5(2+3x)-2(y-2)=0 10+15x-2y+4=0 15x-2y=-14 Above is an \(example\) -------------------------